# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 4 Classical Equivalence of Field Equations Across Formulations
We have seen that the equations of motion derived in terms of the old fields $\mathcal{D}_{ab}$, $\Pi^{a}{}_{bc}$, and $g^{mn}$ are equivalent to those derived in terms of the new fields ${\bar{{\mathcal{P}}}}_{ab}$, ${\tilde{C}}^{a}{}_{bc}$, and $g^{mn}$ upon enforcement of the equations of motion (on-shell). They describe the same classical system, regardless of field choice, as expected. Eq. (3.61) is a particularly interesting relationship between $\tilde{\mathbb{E}}_{c}{}^{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$ as just the right combination to remove the non-tensorial part of ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ to create the tensorial $\tilde{\mathbb{E}}_{c}{}^{ab}$. Furthermore, we see that the tensorial result $\tilde{\mathbb{E}}_{c}{}^{ab}$ is also manifestly projectively invariant as it is composed entirely of such objects ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$.
@@
In the following, we investigate some nuances in the variation of the Lagrangian and summarize the field equations and their classical equivalence. We start by investigating the tensorial or non-tensorial character of the fields and their variations. We suspect that this will have interesting consequences at the quantum level in terms of Faddeev-Popov-like projective gauge fixing conditions. It is therefore important to work out the details at the classical level here first, leaving the quantum calculation for a future endeavor.
4.1 Equivalence of the Variation of the Action
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}\left(\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\Pi^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}\right)~{}~{}~{}.$ (4.1)
Using Eqs. (3.11), (3.1), and (3.8) we can solve for the old field variations $\delta{\mathcal{D}}_{ab}$ and $\delta\Pi^{c}{}_{ab}$ in terms of the new field variations $\delta{\bar{{\mathcal{P}}}}_{ab}$ and $\delta{\tilde{C}}^{c}{}_{ab}$
$\delta\mathcal{D}_{ab}=$ $\delta\bar{\mathcal{P}}_{ab}-\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}-g_{c}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})-g_{c}\delta{\tilde{C}}^{c}{}_{ab}$ (4.2)
$\delta\Pi^{c}{}_{ab}=$ $\delta\hat{\Gamma}^{c}{}_{ab}+\delta{\tilde{C}}^{c}{}_{ab}+\delta^{c}{}_{(a}\delta g_{b)}~{}~{}~{}.$ (4.3)
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}-g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.4)
$-\tfrac{1}{{\rm d}+1}\mathbb{D}^{ab}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}-g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}\big{)}.$ (4.5)
@@
$g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}=$ $(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta{\tilde{C}}^{c}{}_{ab}$ (4.6)
$g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})=$ $(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta\hat{\Gamma}^{c}{}_{ab}~{}~{}~{}.$ (4.7)
@@
Substituting Eqs. (4.6) and (4.7) into Eq. (4.4), the second and third terms on the first line can be combined using Eq.  (3.61), as can the second and third terms on the second line. Furthermore, in Eq. (4.4) the first term on the second line can be integrated by parts. All this leaves us with
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.8)
@@
$+\tfrac{1}{{\rm d}+1}({\bar{\nabla}}_{a}\mathbb{D}^{ab})\delta\hat{\Gamma}^{m}{}_{bm}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}\big{)}~{}~{}~{}.$ (4.9)
Using Eqs. (3.65) and (3.66) we can rewrite the second line in terms of the variation of $\delta g^{mn}$
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{[}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.10)
@@
$+\left(\tfrac{1}{2({\rm d}+1)}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+\tfrac{1}{2}\hat{\nabla}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\delta g^{mn}\right)\big{]}~{}.$ (4.11)
@@
Switching from ^$\hat{\nabla}_{a}$ to ¯${\bar{\nabla}}_{a}$ via Eq. (3.8), we find that the resulting coefficients of all $\delta g^{mn}$ terms combine to form $\mathbb{H}_{mn}$ as defined in Eq. (3.85). This leaves us with our final result:
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}\left(\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathbb{H}_{mn}\delta g^{mn}\right)~{}~{}~{}.$ (4.12)
@@
All individual equations of motion and variations in Eq. (4.12) in terms of the new fields are tensorial. This is not true for Eq. (4.1) in terms of the old fields. It must be true, however, that the variation of the action is general coordinate invariant either way. What saves Eq. (4.1), maintaining its general coordinate invariance, is that the non-tensorial parts of $\delta{\mathcal{D}}_{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ conspire to cancel in Eq. (4.1) while all other fields and variations are tensorial. Specifically, performing a coordinate transformation on Eqs. (4.2) and (2.78) we find
$\delta{\mathcal{D}}^{\prime}_{ab}=$ $\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\delta{\mathcal{D}}_{mn}+j_{c}\delta\Pi^{c}{}_{mn}\right]$ (4.13)
@@
~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}=$ $\frac{\partial x^{p}}{\partial x^{\prime c}}\frac{\partial x^{\prime a}}{\partial x^{\prime m}}\frac{\partial x^{\prime b}}{\partial x^{n}}\left[\tilde{\mathcal{E}}_{p}{}^{mn}-j_{p}\mathbb{D}^{mn}+\tfrac{1}{{\rm d}+1}\delta_{p}{}^{(m}j_{d}\mathbb{D}^{n)d})\ \right]$ (4.14)
$j_{d}=$ $\tfrac{1}{{\rm d}+1}\partial_{m}\log J=\tfrac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{m}\partial x^{\prime q}}~{}~{}~{},~{}~{}~{}J=|\partial x^{p}/\partial x^{\prime q}|$ (4.15)
@@
where the non-tensorial parts of these transformations stem from their dependence on $g_{a}$ as all other parts are tensorial (though a connection itself is not tensorial, its variation is tensorial). Note that the above result for $\delta\mathcal{D}^{\prime}_{ab}$ is consistent with taking the functional variation of Eq. (2.11) as $\delta j_{m}=0$ necessarily.
@@
With the above results for $\delta{\mathcal{D}}^{\prime}_{ab}$ and ~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}$, we find that the following combination found in Eq. (4.1) indeed transforms as a scalar
$\mathbb{D}^{\prime ab}\delta{\mathcal{D}}^{\prime}_{ab}+\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}\delta\Pi^{\prime c}{}_{ab}=\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\Pi^{c}{}_{ab}$ (4.16)
@@
4.2 Summary of Field Equations for Each Formulation
@@
$\mathbb{D}^{ab}=$ $-\frac{d-1}{4J_{0}c\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}=0$ (4.17)
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}=0$ (4.18)
$\mathbb{H}_{mn}=$ $\mathcal{H}_{mn}+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (4.19)
$+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]$ (4.20)
$=$ 00 0 (4.21)
@@
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (4.22)
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (4.23)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (4.24)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (4.25)
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})~{}~{}~{}.$ (4.26)
@@
$\mathbb{D}^{ab}=$ $0\displaystyle 0~{}~{}~{},~{}~{}~{}\mathcal{H}_{mn}=$ 0 (4.27)
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}+\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)=0~{}~{}~{}.$ (4.28)
@@
On shell, these two sets of equations simultaneously reduce to
[…]
@@
Finally, it is useful to express the equations of motion for $g^{mn}$ in the form of Einstein equations. On-shell, this becomes for either formulation
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (4.30)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\left[{\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}\right.$ (4.31)
$\qquad\hskip 10.0pt\left.-\tfrac{1}{2}g_{mn}\left(g^{ab}{\bar{\nabla}}_{c}{\tilde{C}}^{c}{}_{ab}+{\tilde{C}}^{ca}{}_{b}{\tilde{C}}^{b}{}_{ca}+({\rm d}-1){\bar{{\mathcal{P}}}}\right)\right]$

### v3

@@ IV Classical Equivalence of Field Equations Across Formulations
We have seen that the equations of motion derived in terms of the old fields $\mathcal{D}_{ab}$, $\Pi^{a}{}_{bc}$, and $g^{mn}$ are equivalent to those derived in terms of the new fields ${\bar{{\mathcal{P}}}}_{ab}$, ${\tilde{C}}^{a}{}_{bc}$, and $g^{mn}$ upon enforcement of the equations of motion (going on-shell). They describe the same classical system, regardless of field choice, as expected. Equation (178) is a particularly interesting relationship between $\tilde{\mathbb{E}}_{c}{}^{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$ as just the right combination to remove the non-tensorial part of ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ to create the tensorial $\tilde{\mathbb{E}}_{c}{}^{ab}$. Furthermore, we see that the tensorial result $\tilde{\mathbb{E}}_{c}{}^{ab}$ is also manifestly projective invariant as it is composed entirely of such objects ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$.
@@
In the following, we investigate some nuances in the variation of the Lagrangian and summarize the field equations and their classical equivalence across formulations. We start by investigating the tensorial or non-tensorial character of the fields and their variations. We suspect that this will have interesting consequences at the quantum level in terms of Faddeev-Popov-like projective gauge fixing conditions. It is therefore important to work out the details at the classical level here first, leaving the quantum calculation for a future endeavor.
IV.1 Equivalence of the Variation of the Action
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{|g|}\left[\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}^{\ ab}\delta\Pi^{c}_{\ ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}\right].$ (207)
Using Eqs. (112), (104), and (108) we can solve for the old field variations $\delta{\mathcal{D}}_{ab}$ and $\delta\Pi^{c}{}_{ab}$ in terms of the new field variations $\delta{\bar{{\mathcal{P}}}}_{ab}$ and $\delta{\tilde{C}}^{c}{}_{ab}$
$\delta\mathcal{D}_{ab}=$ $\delta\bar{\mathcal{P}}_{ab}-\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}$ (208)
$-g_{c}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})-g_{c}\delta{\tilde{C}}^{c}{}_{ab}$ (209)
$\delta\Pi^{c}{}_{ab}=$ $\delta\hat{\Gamma}^{c}{}_{ab}+\delta{\tilde{C}}^{c}{}_{ab}+\delta^{c}{}_{(a}\delta g_{b)}~{}.$ (210)
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{|g|}\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}$ (211)
$-g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}-\tfrac{1}{{\rm d}+1}\mathbb{D}^{ab}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}$ (212)
$-g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}\big{)}.$ (213)
@@
$g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}=(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta{\tilde{C}}^{c}{}_{ab}$ (214)
$g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})$ (215)
$\qquad\qquad\quad=(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta\hat{\Gamma}^{c}{}_{ab}~{}.$ (216)
@@
Substituting Eqs. (214) and (215) into Eq. (211), the second and third terms on the first line can be combined using Eq.  (178), as can the second and third terms on the second line. Furthermore, in Eq. (211) the first term on the second line can be integrated by parts. All this leaves us with
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{|g|}\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}^{\ ab}\delta{\tilde{C}}^{c}_{\ ab}$ (217)
@@
$+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}+\tfrac{1}{{\rm d}+1}({\bar{\nabla}}_{a}\mathbb{D}^{ab})\delta\hat{\Gamma}^{m}_{\ bm}+\tilde{\mathbb{E}}_{c}^{\ ab}\delta\hat{\Gamma}^{c}_{\ ab}\big{)}.~{}~{}~{}~{}$ (218)
Using Eqs. (185) and (187) we can rewrite the second line in terms of the variation of $\delta g^{mn}$
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{|g|}\Big{[}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}^{\ ab}\delta{\tilde{C}}^{c}_{\ ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$
@@
$\qquad+\left(\tfrac{1}{2({\rm d}+1)}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+\tfrac{1}{2}\hat{\nabla}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}\right)\delta g^{mn}\Big{]}.$ (219)
@@
Switching from ^$\hat{\nabla}_{a}$ to ¯${\bar{\nabla}}_{a}$ via Eq. (108), we find that the resulting coefficients of all $\delta g^{mn}$ terms combine to form $\mathbb{H}_{mn}$ as defined in Eq. (198). This leaves us with our final result:
@@
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{|g|}\left[\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}^{\ ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathbb{H}_{mn}\delta g^{mn}\right].$ (220)
@@
All individual equations of motion and variations in Eq. (220) in terms of the new fields are tensorial. This is not true for Eq. (207) in terms of the old fields. It must be true, however, that the variation of the action is general coordinate invariant either way. What saves Eq. (207), maintaining its general coordinate invariance, is that the non-tensorial parts of $\delta{\mathcal{D}}_{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ conspire to cancel in Eq. (207) while all other fields and variations are tensorial. Specifically, performing a coordinate transformation on Eqs. (208) and (92) we find
$\delta{\mathcal{D}}^{\prime}_{ab}=\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\delta{\mathcal{D}}_{mn}+j_{c}\delta\Pi^{c}{}_{mn}\right]$ (221)
@@
~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}=\frac{\partial x^{p}}{\partial x^{\prime c}}\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial x^{\prime b}}{\partial x^{n}}\left[\tilde{\mathcal{E}}_{p}^{\ mn}-j_{p}\mathbb{D}^{mn}+\tfrac{1}{{\rm d}+1}\delta_{p}^{\ (m}j_{d}\mathbb{D}^{n)d}\right]$ (222)
$j_{c}=\tfrac{1}{{\rm d}+1}\partial_{c}\log J=\tfrac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{c}\partial x^{\prime q}}$ (223)
$J=|\partial x^{p}/\partial x^{\prime q}|$ (224)
@@
where the non-tensorial parts of these transformations stem from their dependence on $g_{a}$ as all other parts are tensorial (though a connection itself is not tensorial, its variation is tensorial). Note that the above result for $\delta\mathcal{D}^{\prime}_{ab}$ is consistent with taking the functional variation of Eq. (13) as $\delta j_{m}=0$ necessarily.
@@
With the above results for $\delta{\mathcal{D}}^{\prime}_{ab}$ and ~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}$, we find that the following combination found in Eq. (207) indeed transforms as a scalar
$\mathbb{D}^{\prime ab}\delta{\mathcal{D}}^{\prime}_{ab}+\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}\delta\Pi^{\prime c}{}_{ab}=\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\Pi^{c}{}_{ab}$ (225)
@@
IV.2 Summary of Field Equations for Each Formulation
@@
$0=\mathbb{D}^{ab}=\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}-({\rm d}-1)\mathbb{K}^{ab}-2\mathcal{K}^{ab}$ (226)
$0=\tilde{\mathbb{E}}_{c}{}^{ab}\equiv\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}$ (227)
$0=\mathbb{H}_{mn}=\mathcal{H}_{mn}+\tfrac{2J_{0}\kappa_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (228)
$\qquad\qquad\hskip 5.0pt+2J_{0}\kappa_{0}\left[{\bar{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}+{\tilde{C}}^{p}_{\ a(m}\tilde{\mathbb{E}}^{*\ \ a}_{n)p}\right]~{}~{}~{}~{}~{}~{}~{}~{}$ (229)
@@
$\mathbb{E}_{c}^{\ ab}={\bar{\nabla}}_{c}\mathbb{K}^{ab}-{\bar{\nabla}}_{d}{\mathcal{K}}_{c}^{\ (ab)d}-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (230)
$\mathbb{K}^{ab}\equiv g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4\mathcal{K}^{ab}$ (231)
$\tilde{\mathbb{E}}^{*\ \ a}_{mn}\equiv\tilde{\mathbb{E}}_{(mn)}^{\ \ \ \ \ a}-\tilde{\mathbb{E}}^{a}_{\ mn}$ (232)
$\mathcal{H}_{mn}=-\mathcal{K}_{mn}+\kappa_{0}\left(2\mathbb{L}_{mn}-g_{mn}\mathcal{L}\right)$ (233)
$\qquad\hskip 12.0pt-g_{mn}\tfrac{2J_{0}\kappa_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (234)
$\mathcal{L}=-\frac{1}{2\kappa_{0}}(\mathcal{K}+2\Lambda_{0})+\mathbb{L}$ (235)
$\mathbb{L}=J_{0}\left[\lambda_{0}^{2}K_{bcd}K^{bcd}-\mathcal{K}_{abcd}\mathcal{K}^{abcd}+4\mathcal{K}_{bd}\mathcal{K}^{bd}-\mathcal{K}^{2}\right]~{}~{}~{}$ (236)
$\mathbb{L}_{mn}=J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}^{\ cd}+2K_{bcm}K^{bc}_{\ \ n})$ (237)
$\qquad\quad+J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}^{\ bcd}-{\mathcal{K}}^{a}_{\ mcd}{\mathcal{K}}_{an}^{\ \ cd}-2{\mathcal{K}}^{abc}_{\ \ \ m}{\mathcal{K}}_{abcn})$
$\qquad\quad+J_{0}\left(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}_{\ n}-2{\mathcal{K}}{\mathcal{K}}_{mn}\right).$ (238)
@@
$\mathbb{D}^{ab}=$ $0\displaystyle 0~{}~{}~{},~{}~{}~{}\mathcal{H}_{mn}=$ 0 (239)
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}+\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)=0~{}.$ (240)
@@
Upon enforcing the equations of motion (going on-shell), these two sets of equations simultaneously reduce to
$\mathbb{D}^{ab}=$ $0~{},~{}\tilde{\mathcal{E}}_{c}^{\ ab}=\tilde{\mathbb{E}}_{c}^{\ ab}=0~{},~{}\mathbb{H}_{mn}=\mathcal{H}_{mn}=0~{}.$ (241)
@@
Finally, it is useful to express the equations of motion for $g^{mn}$ in the form of Einstein equations in a covariant and manifestly projective invariant form. On-shell, this becomes for either formulation
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}+g_{mn}\Lambda_{0}$ (242)
$\hat{T}_{mn}=-\kappa_{0}^{-1}\left[\tilde{Q}_{mn}-\tfrac{1}{2}g_{mn}\tilde{Q}+({\rm d}-1)\left({\bar{{\mathcal{P}}}}_{mn}-\tfrac{1}{2}g_{mn}{\bar{{\mathcal{P}}}}\right)\right]$